Stock XYZ has a current dividend of $3.00 .

The dividend is expected to grow forever at a rate of -1.00% .

Based on the riskiness of XYZ, its discount rate is 4.00%

With this information, what is the dividend yield from today to period 1?

5.00%

-1.10%

-1.00%

5.50%
Answer #1

Answer: Correct answer is 5%
Price of the stock=(Expected year end dividend)/(Required rate of return-Growth rate)
Expected dividend=(Current dividend)*(1+growth rate)=3*(1-1%)
=3*(1-1%)/(4%-(-1%))
=59.4

Dividend yield=Expected dividend/Current price
Expected dividend=(Current dividend)*(1+growth rate)=3*(1-1%)
Dividend yield=3*(1-1%)/59.4=0.05 or 5%